OpenStudy (goten77):

calculus help... really hard integral i think its actually impossible if not then this has to be a joke... question is below

OpenStudy (anonymous):

calculus 1?

OpenStudy (goten77):

\[\int\limits_{?}^{?}\frac{ 4e ^{sinx }+\ln 3x }{ \sqrt{tanx} - \left| 2x \right|}^{2}\]

OpenStudy (goten77):

ignore the question marks on the integral symbol

OpenStudy (anonymous):

I would suggest by parts integration, I know it is a pain but take your time and make sure you are very careful

OpenStudy (goten77):

this was a bonus question on today's test so the answer really doesnt matter but it b intresting to c how its done.. and if by parts man that would b like really long